In a 5.00 km race, one runner runs at a steady 11.5km/h and another runs at 14.0 km/h
How long does the faster runner have to wait at the finish line to see the slower runner cross?
1
Expert's answer
2021-09-03T08:25:19-0400

Given:

"d=5.00\:\rm km"

"v_1=11.5\:\rm km/h"

"v_2=14.0\:\rm km/h"


The time of motion of runners

"t_1=\frac{d}{v_1}=\frac{5.00}{11.5}=0.435\:\rm h"

"t_2=\frac{d}{v_2}=\frac{5.00}{14.0}=0.357\:\rm h"

The difference of time

"\Delta t=t_1-t_2=0.435-0.357=0.078\:\rm h=4.68\: min"